OpenStudy (anonymous):

The area of a rectangular swimming pool is 10x2 – 19x – 15. The length of the pool is 5x + 3. What is the width of the pool?

OpenStudy (anonymous):

Area = length * width. so 10x*2 - 19x - 15 = 5x + 3 * w

OpenStudy (anonymous):

or 10x*2 - 19x - 15 = 5x + 3 * x Simplifying 10x * 2 + -19x + -15 = 5x + 3x Reorder the terms for easier multiplication: 10 * 2x + -19x + -15 = 5x + 3x Multiply 10 * 2 20x + -19x + -15 = 5x + 3x Reorder the terms: -15 + 20x + -19x = 5x + 3x Combine like terms: 20x + -19x = 1x -15 + 1x = 5x + 3x Combine like terms: 5x + 3x = 8x -15 + 1x = 8x Solving -15 + 1x = 8x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-8x' to each side of the equation. -15 + 1x + -8x = 8x + -8x Combine like terms: 1x + -8x = -7x -15 + -7x = 8x + -8x Combine like terms: 8x + -8x = 0 -15 + -7x = 0 Add '15' to each side of the equation. -15 + 15 + -7x = 0 + 15 Combine like terms: -15 + 15 = 0 0 + -7x = 0 + 15 -7x = 0 + 15 Combine like terms: 0 + 15 = 15 -7x = 15 Divide each side by '-7'. x = -2.142857143 Simplifying x = -2.142857143

OpenStudy (precal):

drkole's set up was correct but the work is incorrect. the missing width is (2x-5) because length times width = Area (5x+3)(2x-5)=10x^2-19x-15

OpenStudy (anonymous):

Thank you!

OpenStudy (precal):

Anytime :)